OpenStudy (anonymous):

solve the system of equation by graphing x+y=5 x-y=-3

OpenStudy (anonymous):

y=5-x y=3+x graph that

OpenStudy (anonymous):

so it would be 4 and 4right?

OpenStudy (anonymous):

how though?